# Work statistics for Quantum Spin Chains: characterizing quantum phase transitions, benchmarking time evolution, and examining passivity of quantum states

### v3

@@ III.3 Thermal density matrix as MPO
Instead of adopting METTS, a more direct way of constructing a thermal state based on MPS is to represent the density operator $e^{-\beta H}$ by an MPO. The partition function for evaluating free energy is to take a trace of this MPO. To construct such an MPO, one first prepares an initial identity MPO of bond dimension $\chi_{cut}=1$ to make such an MPO. The MPO for the thermal state is then obtained by acting with the operator $e^{-\beta H}$, which can be decomposed by TEBD and Suzuki-Trotter expansion into a sequence of a product of two-body gates, as shown in Fig. 1 (c). The partition function can be obtained by tracing out the two physical indices of the MPO $e^{-\beta H}$ as shown in Fig. 1(d). In Appendix A.1, we show the agreement on the work statistics obtained from METTS and MPO representation of thermal states. In the rest of the paper, we will adopt MPO to represent thermal states to proceed with the numerical calculations.
@@ IV.1 Quantum spin-$1/2$ chain in magnetic field
where $S_{k}^{x,y,z}$ are the site spin-$1/2$ operators, $k$ labels the number of sites. When $\gamma=1$, it reduces to the quantum Ising chain, and we will focus on the parameter range: $0\leq\gamma\leq 1$ and $h_{x}=0$. The ground state of this model can be exactly solved and exhibits three phases of the Landau-Ginzburg-Wilson paradigm: the oscillatory phase $(O)$, the ferromagnetic phase $(F)$, and the paramagnetic phase $(P)$ [47]. The ground state stays in the $O$ phase for small $h_{z}$ until $h_{z}={1\over 2}\sqrt{1-\gamma^{2}}$ and then changes to the $F$ phase as a crossover transition. Further increasing $h_{z}$ up to $h_{z}=0.5$, the ground state will change to $P$ phase, which is a second-order quantum phase transition in the thermodynamic limit with $\langle S_{x}\rangle$ as the order parameter, i.e., $\langle S_{x}\rangle$ is nonzero in $F$ (or $O$ if $\gamma=0$) phase but is zero in $P$ phase. Below, we will consider the phase transition between the ferromagnetic and paramagnetic phases.
Figure 2: Work characteristic function $G(-is)$ under a sudden quench process implemented by Hamiltonian $H=H_{0}+\Delta\lambda\;V$ with $\Delta\lambda=0.01$ with $H_{0}$ of a $n=100$ quantum Ising chain, i.e., (31) of $\gamma=1$, $h_{x}=0$ and $h_{z}\in[0,2]$, and $V=\sum_{k}S_{k}^{z}$. The results are obtained by the MPS-based TEBD method via the 1-st order Trotter expansions for $\delta=0.01$ with bond dimension $\chi_{cut}=16$. We show the results only for five values of $s$ but can connect them into a line by Lagrange interpolation. Each color represents one particular choice of $h_{z}$, with their values indicated by the color sidebar. The Lagrange interpolating lines at critical points are shown explicitly for $h_{z}=0.5$. The slope of an interpolating line corresponds to the average work done $\overline{W}$, which is mostly negative.
Figure 3: Phase diagrams of quantum Heisenberg XY spin-$1/2$ chain as functions of $h_{z}$ at zero temperature of size $n=100$ for $\gamma=0.5$ (red crosses), $0.8$ (blue circles) and $1.0$ (green squares). They are respectively characterized by (a) entanglement entropy, (b) the average work done per site, $\overline{w}$, under a sudden quench process with $V=\sum_{k}S_{k}^{z}$ and $\Delta\lambda=0.01$, and its associated higher moments and cumulants: (c) the second moment of $\overline{w^{2}}$, (d) the second cumulant $\sigma^{2}_{W}=\kappa_{2}=\overline{w^{2}}-\overline{w}^{2}$, (e) the third cumulant $\kappa_{3}=\overline{w^{3}}-3\overline{w^{2}}\cdot\overline{w}+\overline{w}^{3}$, and (f) the fourth cumulant $\kappa_{4}$. The plots are also obtained by the TEBD method with the parameters $\chi_{cut}$ and $\delta$ shown on the subfigures. As expected, the higher moments or cumulants yield sharper change near the critical point, so that the work statistics can indicate the quantum phase transitions in the Ising-like spin chains.
@@
Here, we apply the MPS-based TEBD method to calculate the characteristic function $G(u)$ of (10) for this spin chain model with $\gamma=1$ and $h_{x}=0$ at zero temperature but varying the transverse field $h_{z}$. We aim to check if the work done $\overline{W}$ can be used as an order parameter for the quantum phase transition, i.e., from $F$ (or $O$ if $\gamma=0$) phase to $P$ phase. For simplicity, we consider the sudden quench process implemented by the Hamiltonian of (11) and (12) with a chosen jump of the coupling constant, i.e., $\Delta\lambda=0.01$. The results are shown in Fig. 2 and 3.
@@
The Fig. 2 shows the characteristic function $G(-is)$ of a $n=100$ Ising chain for $s\in[-0.2,0.2]$ with $\gamma=1$ and $h_{x}=0$ at various values of $h_{z}$ in the initial Hamiltonian $H_{0}$. For simplicity, we pick five values of $s$ but vary $h_{z}$ continuously with its value indicated by the color bar attached aside. In particular, the five points belonging to $h_{z}=0.5$ are denoted by the crosses and joined into a line by Lagrange interpolation. We see that the slope of this line is negative, which implies that the average work done $\overline{W}=\lim_{s\rightarrow 0}{\partial G(-is)\over\partial s}$ is negative. For other values of $h_{z}$ in the initial Hamiltonian $H_{0}$, we can see that the slope of the line connected by the five points of the same $h_{z}$ starts with a positive value for $h_{z}=0$, then gradually turns negative as $h_{z}$ increases and finally converges to a constant negative value.
@@
The above results and further related calculations can be translated into phase diagrams as shown in Fig. 3 for the quantum phase transition from the $F$ (or $O$) phase to the $P$ phase. These phase diagrams are characterized by the entanglement entropy $S$, the average work done per site, $\overline{w}$, and its associated higher moments and cumulants, under a sudden quench process with $V=\sum_{k}S_{k}^{z}$ and $\Delta\lambda=0.01$. We see that $S$ has a sharp peak near $h_{z}=0.5$, which implies a stronger correlation near the critical point at $h_{z}=0.5$. Similarly, the work statistics denoted by $\overline{w}$ and the higher moments and cumulants also show the shaper changes near $h_{z}=0.5$. As expected, the higher moments and cumulants show sharper changes at the critical point. This implies that the work statistics can indicate the phase transitions. Interestingly, the cumulants of even orders show a level-off for $h_{z}\leq 0.5$.
@@
Finally, in Appendix A.2, we give a consistency check to show that the above numerical results of $\overline{W}$ from either Eq. (7) or Eq. (16) agree.
@@ IV.2 Anisotropic quantum spin-$1$ chain
Figure 4: Phase diagrams of quantum Heisenberg XXZ spin-$1$ chain as a function of coupling parameter $D$ at zero temperature of lattice size $n=100$ for $J_{z}=1.0$ (red crosses), $1.5$ (blue circles). As in Fug. 3, they are characterized by the entanglement entropy $S$ to indicate the quantum critical points with sharp peaks, and by (b) the average work done per site $\overline{w}$ under a quenching process with $V=\sum_{k}(S_{k}^{z})^{2}$ and $\Delta\lambda=0.01$, and by its associated (c) second moment and (d) second cumulant. The plots are also obtained by the TEBD method with the parameters $\chi_{cut}$ and $\delta$ shown on the subfigures. Unlike in Fig. 3, we see that $\overline{w}$ and $\overline{w^{2}}$ cannot capture the critical points at $D=1.0$ for $J_{z}=1.0$ and $D=0.5$ and $D=1.34$ for $J_{z}=1.5$ as indicated by the peaks of $S$. However, the second cumulant $\sigma^{2}_{w}$ does show some crossover behavior near these critical points. Overall, the work statistic can indicate the quantum phase transitions in the Haldane-like spin chains.
As before, we can characterize the quantum phase transitions of this model by entanglement entropy $S$ and the average work done per site $\overline{w}$ by a sudden quench with $V=\sum_{k}(S_{k}^{z})^{2}$ and $\Delta\lambda=0.01$, and by the associated higher moments and cumulants. The results are shown in Fig. 4, which are again obtained numerically using the MPS-based TEBD method. We see that the entanglement entropy $S$ shows the peaks around the quantum critical points, which are at $D=1.0$ for $J_{z}=1.0$ and $D=0.5$ and $D=1.34$ for $J_{z}=1.5$. Instead, $\overline{w}$ and the second moment $\overline{w^{2}}$ behaves smoothly. Despite that, the second cumulant $\sigma^{2}_{w}$ does show some crossover behavior near these critical points. From this trend and the similar one shown in Fig.  3, we expect the higher moments and cumulants to display shaper behavior around the quantum critical points, although it is quite computationally costly to obtain higher moments and cumulants due to the need of large bond dimension. Thus, we again can conclude that the work statistics can indicate quantum phase transitions in the Haldane-like spin chains.
@@ V Benchmarking the real-time evolving methods by Jarzynski’s equality
We now adopt the Jarzynski equality of (9) dictated by the real-time correlator expression of the work characteristic function $G(i\beta;t_{f})|_{\beta=1}$, i.e., (10), to benchmark the numerical accuracy of the real-time evolution of two quantum spin chain models considered above. The numerical error mainly comes from evaluating $G(i\beta;t_{f})$ and accumulates as $t_{f}$ grows. For the small-size cases, we can use either the Exact Diagonalization (ED) method or the MPS-based TEBD method. For large-size cases, we can only use the TEBD method. As before, the non-equilibrium process is implemented by the Hamiltonian of (11), and here we adopt the linear profile of coupling constant, i.e., $\lambda(t)=t$. The results for the small-size cases are shown in Fig. 5 for (a) $n=10$ Ising chain of fixed $h_{x,z}$ and with $V=\sum_{i}S_{i}^{z}$ and (b) $n=6$ Haldane chain of fixed $D=2$ and with $V=\sum_{i}(S_{i}^{z})^{2}$. Fig. 5 shows that as $t_{f}$ increases, the curve of $G(i\beta;t_{f})|_{\beta=1}$ of the Ising chain increases; on the other hand, the curve in the Haldane chain decreases.
Figure 6: Benchmarking the ED and TEBD methods by the ratio $R(t_{f})={G(i\beta;t_{f})\over Z(t_{f})/Z(0)}\Big{|}_{\beta=1}$ from Jarzynski equality for subfigure (a) $n=10$ Ising chain with $h_{x}=h_{z}=1$ and $V=\sum_{k}S_{k}^{z}$; and subfigure (b) $n=6$ Haldane chain with $D=2$ and $V=\sum_{k}(S_{k}^{z})^{2}$. In subfigure (c), we show the dependence of the resultant $R(t_{f})$ of $n=6$ Haldane chain on the bond dimension $\chi_{cut}$ of MPS. As expected, larger $\chi_{cut}$ yields better results as expected. The results from ED are indicated by squares and from TEBD by circles. Besides, we also show the ratio ${Z(t_{f})\over Z(0)}|_{\rm MPS}\Big{/}{Z(t_{f})\over Z(0)}|_{\rm ED}$ (triangles in subfigures (a) and (b)) to demonstrate the accuracy of TEBD in evaluating the partition functions. In all the above, the coupling constant has the linear profile, i.e., $\lambda(t)=t$. If there is no numerical error, $R=1$, and the deviation benchmarks the real-time numerical errors. For ED, we see that $R$ remains equal to one. On the other hand, for MPS, we see that the real-time numerical errors start to escalate around $t_{f}=0.7$. However, the ratio of partition functions remains accurate.
@@
Next, we use Jarzynski equality to characterize the numerical error of real-time evolution and benchmark the corresponding numerical methods. This can be quantified by the ratio $R(t_{f})={G(i\beta;t_{f})\over Z(t_{f})/Z(0)}|_{\beta=1}$ defined in (14), which is the ratio between either side of Jarzynski equality. If there is no numerical error, the Jarzynski equality is obeyed so that $R=1$. Otherwise, the $|1-R|$ size can characterize the numerical error.

### v4

@@ III.3 Thermal density matrix as MPO
Instead of adopting METTS, a more direct way of constructing a thermal state based on MPS is to represent the density operator $e^{-\beta H}$ by an MPO. The partition function for evaluating free energy is to take a trace of this MPO. To construct such an MPO, one first prepares an initial identity MPO of bond dimension $\chi_{cut}=1$ to make such an MPO. The MPO for the thermal state is then obtained by acting with the operator $e^{-\beta H}$, which can be decomposed by TEBD and Suzuki-Trotter expansion into a sequence of a product of two-body gates, as shown in Fig. 1 (c). The partition function can be obtained by tracing out the two physical indices of the MPO $e^{-\beta H}$ as shown in Fig. 1(d). In Appendix B.1, we show the agreement on the work statistics obtained from METTS and MPO representation of thermal states. In the rest of the paper, we will adopt MPO to represent thermal states to proceed with the numerical calculations.
Figure 2: Work characteristic function $G(-is)$ under a sudden quench process implemented by Hamiltonian $H=H_{0}+\Delta\lambda\;V$ with $\Delta\lambda=0.01$ with $H_{0}$ of a $n=1024$ quantum Ising chain, i.e., (33) of $\gamma=1$, $h_{x}=0$ and $h_{z}\in[0,2]$, and $V=\sum_{k}S_{k}^{z}$. The results are obtained by the MPS-based TEBD method via the 1-st order Trotter expansions for $\delta=0.01$ with bond dimension $\chi_{cut}=16$. We show the results only for five values of $s$ but can connect them into a line by Lagrange interpolation. Each color represents one particular choice of $h_{z}$, with their values indicated by the color sidebar. The Lagrange interpolating lines at critical points are shown explicitly for $h_{z}=0.5$. The slope of an interpolating line corresponds to the average work done $\overline{W}$, which is mostly negative.
@@ IV.1 Quantum spin-$1/2$ chain in magnetic field
Figure 3: Moments and cumulants of work done by sudden quench with $V=\sum_{k}S_{k}^{z}$ and $\Delta\lambda=0.01$ on quantum Heisenberg XY Ising chain of size $n=1024$ as a function of the transverse field $h_{z}$ at zero temperature for $\gamma=0.5$ (red crosses), $0.8$ (blue circles) and $1.0$ (green squares). For comparison, the entanglement entropy $S$ is also shown. They are listed as follows: (a) $\overline{W}$, (b) $\overline{W^{2}}$, (c) $\overline{W^{3}}$, (d) $\overline{W^{4}}$, (e) $S$, (f) $\sigma^{2}_{W}$, (g) $\kappa_{3}$ and (h) $\kappa^{4}$. The plots are obtained by the TEBD method with the parameters $\chi_{cut}=16$ and $\delta=0.01$, and $\overline{W^{n}}$ are obtained using Eq. (16) by the MPO method. As expected, the higher moments or cumulants yield sharper change near the critical point so that the work statistics can indicate the quantum phase transitions in the Ising-like spin chains. To go beyond sudden quench and for comparison, we also extract $\overline{W}$ and $\overline{W^{2}}$ (black dashed lines) from the characteristic function Eq. 10 for the non-sudden quench process with $t_{f}=0.1$.
@@
where $S_{k}^{x,y,z}$ are the site spin-$1/2$ operators, $k$ labels the number of sites. When $\gamma=1$, it reduces to the quantum Ising chain, and we will focus on the parameter range: $0\leq\gamma\leq 1$ and $h_{x}=0$. The ground state of this model can be exactly solved and exhibits three phases of the Landau-Ginzburg-Wilson paradigm: the oscillatory phase $(O)$, the ferromagnetic phase $(F)$, and the paramagnetic phase $(P)$ [66]. The ground state stays in the $O$ phase for small $h_{z}$ until $h_{z}={1\over 2}\sqrt{1-\gamma^{2}}$ and then changes to the $F$ phase as a crossover transition. Further increasing $h_{z}$ up to $h_{z}=0.5$, the ground state will change to $P$ phase, which is a second-order quantum phase transition in the thermodynamic limit with $\langle S^{x}\rangle$ as the order parameter, i.e., $\langle S^{x}\rangle$ is nonzero in $F$ (or $O$ if $\gamma=0$) phase but is zero in $P$ phase. Below, we will consider the phase transition between the ferromagnetic and paramagnetic phases.
@@
Here, we apply the MPS-based TEBD method to calculate the characteristic function $G(u)$ of Eq. (10) for this spin chain model with $\gamma=1$ and $h_{x}=0$ at zero temperature but varying the transverse field $h_{z}$. We aim to check if the work done $\overline{W}$ can be used as an order parameter for the quantum phase transition, i.e., from $F$ (or $O$ if $\gamma=0$) phase to $P$ phase. For simplicity, we consider the sudden quench process implemented by the Hamiltonian of Eq. (11) and Eq. (12) with a chosen jump of the coupling constant, i.e., $\Delta\lambda=0.01$. The results are shown in Fig. 2 and 3.
@@
The Fig. 2 shows the characteristic function $G(-is)$ of a $n=1024$ Ising chain for $s\in[-0.2,0.2]$ with $\gamma=1$ and $h_{x}=0$ at various values of $h_{z}$ in the initial Hamiltonian $H_{0}$. For simplicity, we pick five values of $s$ but vary $h_{z}$ continuously with its value indicated by the color bar attached aside. In particular, the five points belonging to $h_{z}=0.5$ are denoted by the crosses and joined into a line by Lagrange interpolation. We see that the slope of this line is negative, which implies that the average work done $\overline{W}=\lim_{s\rightarrow 0}{\partial G(-is)\over\partial s}$ is negative. For other values of $h_{z}$ in the initial Hamiltonian $H_{0}$, we can see that the slope of the line connected by the five points of the same $h_{z}$ starts with a positive value for $h_{z}=0$, then gradually turns negative as $h_{z}$ increases and finally converges to a constant negative value.
@@
The above results and further related calculations can be translated into phase diagrams as shown in Fig. 3 for the quantum phase transition from the $F$ (or $O$) phase to the $P$ phase. These phase diagrams are characterized by the entanglement entropy $S$, the average work done, $\overline{W}$, and its associated higher moments and cumulants, under a sudden quench process with $V=\sum_{k}S_{k}^{z}$ and $\Delta\lambda=0.01$. We see that $S$ has a sharp peak near $h_{z}=0.5$, which implies a stronger correlation near the critical point at $h_{z}=0.5$. Similarly, the work statistics denoted by $\overline{W}$ and the higher moments and cumulants also show the shaper changes near $h_{z}=0.5$. As expected, the higher moments and cumulants show sharper changes at the critical point. This implies that the work statistics can indicate the phase transitions. Interestingly, the cumulants of even orders show a level-off for the $F$ (or $O$) phase with $h_{z}\leq 0.5$, and similar behavior for $\sigma^{2}_{W}$ has been observed earlier obtained by the exact solution of Ising chains [67]. This constancy could be associated with some underlying symmetry. However, the quantities $\langle V^{n}\rangle$ with $V=\sum_{k}S_{k}^{z}$, related to the cumulants of work, are different from the order parameter $\langle\sum_{k}S_{k}^{x}\rangle$ for the transition from $F$ (or $O$) phase to the $P$ phase.
@@
In addition, in Appendix B.2, we give a consistency check to show in Fig. 13 that the above numerical results of $\overline{W}$ and $\overline{W^{2}}$ obtained from either Eq. (7) or Eq. (16) agree.
Figure 4: Moments and cumulants of work done by sudden quench with $V=\sum_{k}(S_{k}^{z})^{2}$ and $\Delta\lambda=0.01$ on quantum Heisenberg XXZ Haldane chain of size $n=1024$ as a function of the coupling parameter $D$ at zero temperature for $J_{Z}=1.0$ (red crosses) and $1.5$ (blue circles). For comparison, the entanglement entropy $S$ is also shown. The listing order for $\overline{W^{m=1,\cdots,4}}$, $\sigma^{2}_{W}$, $\kappa_{\ell=3,4}$ and $S$, the subfigure captions and the numerical implementing methods are the same as the one in Fig. 3. The quantum critical points are indicated by the sharper behavior of $S$, which are at (i) $(J_{z},D)=(1.5,0.48)$, (ii) $(J_{z},D)=(1.5,1.34)$ and (iii) $(J_{z},D)=(1.0,1.0)$. Note that (i) is the transition point from the ordered phase to the topological phase, but (ii) and (iii) are the purely topological ones without any order parameter and should be more difficult to detect. For (i), $\overline{W^{m=1,\cdots,4}}$ show very mild behavior, but cumulants show shaper changes as the order of cumulants goes higher. However, for (ii) and (iii), only $\sigma^{2}_{W}$ and $\kappa_{4}$ show mild crossover (see the enlargement insets in the subfigures (f) and (h)), and all the other moments/cumulants in this figure show no obvious changes. Overall, we can conclude that the work statistics can barely capture the purely topological phase transitions. The plots are obtained by the TEBD method with the parameters $\chi_{cut}=64$ and $\delta=0.01$.
@@ IV.2 Anisotropic quantum spin-$1$ chain
As before, we can characterize the quantum phase transitions of this model by entanglement entropy $S$, the average work done $\overline{W}$ and its higher moments/cumulants by a sudden quench with $V=\sum_{k}(S_{k}^{z})^{2}$ and $\Delta\lambda=0.01$, and by the associated higher moments and cumulants. The results are shown in Fig. 4, which are again obtained numerically using the MPS-based TEBD method. We see that the entanglement entropy $S$ shows the peaks around three quantum critical points, which are at (i) $(J_{z},D)=(1.5,0.48)$, (ii) $(J_{z},D)=(1.5,1.34)$ and (iii) $(J_{z},D)=(1.0,1.0)$. The critical point (i) is the transition point from the ordered Neel phase to the topological Haldane phase and is partly with local order parameters. On the other hand, the critical points (ii) and (iii) are from the Haldane phase to the large $D$ phase, and no local parameter exists to characterize it. Therefore, these transitions are purely topological types and should be more difficult to detect. Indeed, this is what we see from Fig. 4. For (i), $\overline{W^{m=1,\cdots,4}}$ show very mild behavior, but cumulants show shaper changes as the order of cumulants goes higher. However, for (ii) and (iii), only $\sigma^{2}_{W}$ shows mild crossover, and all the other moments/cumulants in this figure show no obvious changes. It is strange why the higher cumulants do not show sharper behavior for critical points (ii) and (iii). This could be due to the need for a higher bond dimension to capture the behavior of higher cumulants. Overall, we can conclude that the work statistics can barely capture the purely topological phase transitions. Since work statistics involve all multipoint correlations, and could be used as some non-local order parameter. Thus, we expect the higher moments and cumulants of the work statistics to display shaper behavior around the topological phase transitions, although it is quite computationally costly to obtain higher moments and cumulants due to the need for large bond dimensions.
Also, in Appendix B.2, we give a consistency check to show in Fig. 14 that the above numerical results of $\overline{W}$ obtained from either Eq. (7) or Eq. (16) agree. However, the $\overline{W^{2}}$ obtained from either Eq. (7) or Eq. (16) can’t match well. The larger bound dimension $\chi_{cut}$ might improve the physical properties near critical points.
@@ V Benchmarking the real-time evolving methods by Jarzynski’s equality
We now adopt the Jarzynski equality of Eq. (9) dictated by the real-time correlator expression of the work characteristic function $G(i\beta;t_{f})|_{\beta=1}$, i.e., Eq. (10), to benchmark the numerical accuracy of the real-time evolution of two quantum spin chain models considered above. The numerical error mainly comes from evaluating $G(i\beta;t_{f})$ and accumulates as $t_{f}$ grows. For the small-size cases, we can use either the Exact Diagonalization (ED) method or the MPS-based TEBD method. For large-size cases, we can only use the TEBD method. As before, the non-equilibrium process is implemented by the Hamiltonian of Eq. (11), and here, we adopt the linear profile of coupling constant, i.e., $\lambda(t)=t$. The results for the small-size cases are shown in Fig. 5 for (a) $n=10$ Ising chain of fixed $h_{x,z}$ and with $V=\sum_{i}S_{i}^{z}$ and (b) $n=6$ Haldane chain of fixed $D=2$ and with $V=\sum_{i}(S_{i}^{z})^{2}$. Fig. 5 shows that as $t_{f}$ increases, the curve of $G(i\beta;t_{f})|_{\beta=1}$ of the Ising chain increases; on the other hand, the curve in the Haldane chain decreases.
Figure 6: Benchmarking the ED and TEBD methods by the ratio $R(t_{f})={G(i\beta;t_{f})\over Z(t_{f})/Z(0)}\Big{|}_{\beta=1}$ from Jarzynski equality for subfigure (a) $n=10$ Ising chain with $h_{x}=h_{z}=1$ and $V=\sum_{k}S_{k}^{z}$; and subfigure (b) $n=6$ Haldane chain with $D=2$ and $V=\sum_{k}(S_{k}^{z})^{2}$. In subfigure (c), we show the dependence of the resultant $R(t_{f})$ of $n=6$ Haldane chain on the bond dimension $\chi_{cut}$ of MPS. As expected, larger $\chi_{cut}$ yields better results as expected. The results from ED are indicated by squares, and from TEBD, they are indicated by circles. Besides, we also show the ratio ${Z(t_{f})\over Z(0)}|_{\rm MPS}\Big{/}{Z(t_{f})\over Z(0)}|_{\rm ED}$ (triangles in subfigures (a) and (b)) to demonstrate the accuracy of TEBD in evaluating the partition functions. In all the above, the coupling constant has the linear profile, i.e., $\lambda(t)=t$. If there is no numerical error, $R=1$, and the deviation benchmarks the real-time numerical errors. For ED, we see that $R$ remains equal to one. On the other hand, for MPS, we see that the real-time numerical errors start to escalate around $t_{f}=0.7$. However, the ratio of partition functions remains accurate.
@@
Next, we use Jarzynski equality to characterize the numerical error of real-time evolution and benchmark the corresponding numerical methods. This can be quantified by the ratio $R(t_{f})={G(i\beta;t_{f})\over Z(t_{f})/Z(0)}|_{\beta=1}$ defined in Eq. (14), which is the ratio between either side of Jarzynski equality. If there is no numerical error, the Jarzynski equality is obeyed so that $R=1$. Otherwise, the $|1-R|$ size can characterize the numerical error.